OpenStudy (anonymous):

x^2+2x-3>0 why is the answer -3 and 1 not 3 and -1

OpenStudy (tkhunny):

x^2 + 2x - 3 = (x+3)(x-1) x+3 = 0 x = -3 x-1 = 0 x = 1

OpenStudy (anonymous):

i get it